OpenStudy (anonymous):

Simplify the expression -12 + 2(3x + 5) + 4x

OpenStudy (anonymous):

1) distribute\[a(x+b)=ax+ab\]2) Combine like terms\[ax + c + bx+d=(a+b)x+(c+d)\]

OpenStudy (waleed_imtiaz):

solve brackets and then add and subtract

OpenStudy (anonymous):

-12 + 2(3x + 5) + 4x (multiply through the parenthesis) -12 + 6x + 10 + 4x (combine like terms) 10x - 2 do you need to factor ? If so... 10x - 2 = 2(5x - 1)

OpenStudy (anonymous):

The answer is 10x - 2 or if you need it factored it is 2(5x - 1)